OpenStudy (anonymous):

find the sum of the first 5 terms of the geometric series -2,-6,-18...

OpenStudy (whpalmer4):

Are you able to find the next two terms? Hint: a geometric sequence means that the ratio between any two adjacent terms is always the same value. -6/-2 = -18/-6 = 4th term / -18 = 5th term / 4th term =

OpenStudy (johnweldon1993):

\[S_n = a \frac{ 1-r^n }{ 1-r }\] n = the number of terms you want the sum of a = the first term in the sequence r = the common ratio The common ratio looks like 3 right? -2 X 3 = -6...and -6 X 3 = -18 so yes 3 is the common ratio "r" n Would be 5 here since you want the sum of the first 5 numbers a would be -2 , the first term of the sequence So \[S_5 = -2 \frac{ 1-(3)^5 }{ 1-3 }\] can you solve this?

OpenStudy (anonymous):

no im lost